# On some aspects of local thermodynamical formalism

### v2

@@ Abstract
In 2007, Ye & Zhang introduced a version of local topological entropy. Since their entropy function is, as we show under mild conditions, constant for topologically transitive dynamical systems, we propose to adjust the notion in a way that does not neglect the initial transient part of an orbit. We investigate the properties of this “transient” version, which we call translocal entropy, and compute it in terms of Lyapunov exponents for various dynamical systems. We also investigate how this adjustment affects measure-theoretic local (Brin-Katok) entropy and local pressure functions, generalizing some partial variation principles of Ma & Wen.
Mathematics Subject Classification (2010): Primary: 37B40, Secondary: 37A35,
Keywords. topological entropy, local entropy, Brin-Katok entropy, variational principle, topological pressure, local measure-theoretic pressure
@@ 1 Introduction and motivation
[…]
Local versions of topological entropy were used by Bowen [9] to get an upper estimate of the difference between $h_{\mu}(f)$ and the entropy of a partition with diameter less than $\varepsilon$ and by Misiurewicz [29], who proved that vanishing of his local entropy implies existence of maximal measure. Different versions were used by Newhouse [30], and by and Buzzi & Ruette [13, 15], to address questions about the (unique) existence of a measure of maximal entropy. All these notions rely on Bowen balls (also called dynamical balls):
@@
A local view on measure-theoretic entropy goes back to at least Brin & Katok [11]. In particular, Brin-Katok entropy $h_{\mu}(f,x)$ (see (6) in Section 3.1 for the definition), when integrated over the space yields the global measure-theoretic entropy:
[…]
@@
In 2007, Ye & Zhang [39] introduced a new version of local entropy, which they called entropy function $h_{\text{\tiny top}}:X\to[0,\infty]$. Their definition, see (3) below, is the starting point of our paper, but one soon discovers that under mild conditions the entropy function is constant: $h_{\text{\tiny top}}(z)\equiv h_{\text{\tiny top}}(f)$ for very natural dynamical systems, see Theorem 2.1. It is not straightforward to find maps with non-constant entropy functions, see Section 2.1. The reason is that in the Ye & Zhang approach, only the limit behaviour of the orbit of $z$ counts, not the “transient” behaviour at finite time scales. The entropy function does not see if entropy is created “immediately” in small neighbourhoods of $z$, or only at very distant time scales.
For this reason, we propose an adjustment to Ye & Zhang’s definition, which we call the translocal entropy function and denote as $h_{\omega}$, that takes into account the time scale needed for Bowen balls to reach unit size. As a result, $h_{\omega}(z)$ depends non-trivially on $z\in X$, and Lyapunov exponents at $z$ play a central role. This interaction is most transparent in smooth one-dimensional maps, see Theorem 2.2, where each point has at most one Lyapunov exponent (because there is only one direction in which derivatives can be taken). As representatives of higher-dimensional maps, we study toral automorphisms $F_{A}:{\mathbb{T}}^{d}\to{\mathbb{T}}^{d}$ based on a $d\times d$ matrix $A$. Theorem 2.3 shows that its translocal entropy function is determined by the Lyapunov exponents, i.e., the logarithms of the eigenvalues of $A$.
@@ 1.2 Structure of the paper
Our paper is organized as follows: In Section 2, we introduce the entropy function by Huang & Ye, and study when it is constant and when not. In Section 2.2, we give our adaptation of the entropy function, and give its basic and more advanced properties in the remaining subsections of Section 2.
@@ 2 Local topological entropy in sense of Ye & Zhang
Let $f:X\to X$ be a continuous map on a compact metric space $(X,d)$. For closed subset $K\subset X$ we define the topological entropy $h_{\text{\tiny top}}(f,K)$ restricted to $K$ as the limit of exponential growth rates of $(n,\varepsilon)$-separated subsets of $K$ as $\varepsilon\to 0$, that is
@@
where $S(n,\varepsilon,K)$ is the maximal cardinality of $(n,\varepsilon)$-separated subsets of $K$. A point $z\in X$ is an entropy point if there are arbitrarily small closed neighbourhoods $K$ such that $h_{\text{\tiny top}}(f,K)>0$ and a full entropy point if in addition $h_{\text{\tiny top}}(f,K)=h_{\text{\tiny top}}(f)$, see [34, Definition 5.1] and [39, Section 2]. Hence, only maps with positive entropy can have full entropy points.
It seems natural to define the local entropy of $f$ at a point $z$ as
$\lim_{\delta\to 0}\inf\{h_{\text{\tiny top}}(f,K):K\text{ is a closed $\delta$-neighbourhood of }z\}.$ (2)
However, in [34] this step is not taken, and in fact, if $f$ is locally eventually onto, for instance, then there is $N_{K}$ such that $f^{N_{K}}(K)=X$. Notice that $(n-N_{k},\varepsilon)$-separated points $y_{1},y_{2}\in f^{N_{k}}(K)$ determine points $x_{1},x_{2}\in K$, with $y_{1}=f^{N_{k}}(x_{1})$ and $y_{2}=f^{N_{k}}(x_{2})$, such that $x_{1},x_{2}$ are $(n,\varepsilon)$-separated. Therefore $S(n,\varepsilon,K)\geq S(n-N_{k},\varepsilon,f^{N_{k}}(K))$. Thus the above quantity in (2) is constant $h_{\text{\tiny top}}(f)$.
Instead, in [39, Definition 4.1] the local entropy function is defined as
$h_{\text{\tiny top}}(x)=\lim_{\varepsilon\to 0}h_{\text{\tiny top}}(x,\varepsilon),$ (3)
where
$h_{\text{\tiny top}}(x,\varepsilon)=\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K):K\text{ is a closed $\delta$-neighbourhood of }x\}.$
This differs from (2) by the order in which the limits $\varepsilon\to 0$ and $\delta\to 0$ are taken. In general, $h_{\text{\tiny top}}(x)=h_{\text{\tiny top}}(f(x))$, so the local entropy is constant along forward orbits, see [39, Proposition 4.4(1)]. In fact, we have
@@
Proof. Suppose $x\in X$ has a dense orbit. Let $y\in Y$ with closed neighbourhood $K_{y}$ be arbitrary. Then there is $n\geq 0$ such that $f^{n}(x)\in K_{y}$ and $K_{x}$ such that $f^{n}(K_{y})\subset K_{x}$. It follows that $h_{\text{\tiny top}}(f,K_{x})\leq h_{\text{\tiny top}}(f,K_{y})$, and $h_{\text{\tiny top}}(x)\leq h_{\text{\tiny top}}(y)$ follows. $\Box$
@@ 2.1 When is the entropy function non-constant?
The next theorem show the limitations of the local entropy function: under mild assumptions it is constant. A version of this theorem, for much more specific dynamical systems, can be found in [8].
A map $f:X\to X$ is topologically exact or locally eventually onto if for every open $U\subset X$, there is $N\in{\mathbb{N}}$ such that $f^{N}(U)=X$. Topological exactness gives a simple sufficient condition for the entropy function to be constant.
@@
If $f:X\to X$ is topologically exact, then $h_{\text{\tiny top}}(z)\equiv h_{\text{\tiny top}}(f)$ for any $z\in X$.
@@
Proof. Given a closed neighbourhood $K\owns z$, take $N$ such that $f^{N}(K)\supset f^{N}(\mathring{K})=X$. Then $\frac{1}{n}\log S(n,\varepsilon,K)\geq\frac{1}{n}\log S(n-N,\varepsilon,X)=\frac{n-N}{n}\frac{1}{n-N}\log S(n-N,\varepsilon,X)$. So after taking limits $n\to\infty$ and $\varepsilon\to 0$, we find $h_{\text{\tiny top}}(z)\geq h_{\text{\tiny top}}(f)$. $\Box$
Theorem 2.1
@@
Let $f:X\to X$ be a continuous topological transitive map defined on a compact metric space $(X,d)$. If $h_{\text{\tiny top}}(f)=0$ or if each measure of maximal entropy is fully supported, then $h_{\text{\tiny top}}(z)\equiv h_{\text{\tiny top}}(f)$ for every $z\in X$.
Proof. Since $h_{\text{\tiny top}}(f)\geq h_{\text{\tiny top}}(z)$ for every $z\in X$, only the case that $h_{\text{\tiny top}}(f)>0$ needs a proof.
Assume by contradiction that there is $z\in X$ is such that $h_{\text{\tiny top}}(z)<h_{\text{\tiny top}}(f)$. Take compact neighbourhoods $K^{\prime},K$ of $z$ such that $K^{\prime}$ is compactly contained in $K$, such that
@@
$\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K)<b:=\frac{h_{\text{\tiny top}}(z)+h_{\text{\tiny top}}(f)}{2}.$
The set $Y=\overline{X\setminus\bigcup_{n\geq 0}f^{-n}(K^{\prime})}$ is compact and forward invariant. Since measures of maximal entropy are fully supported, $h^{\prime}:=h_{\text{\tiny top}}(f|_{Y})<h_{\text{\tiny top}}(f)$. By the variational principle, we can take a sequence of $f$-invariant probability measures $\mu_{n}$ such that $h_{\mu_{n}}(f)\to h_{\text{\tiny top}}(f)$. Take $n$ so large that $h_{\mu_{n}}(f)>h^{\prime}$. This means in particular that $\operatorname{supp}(\mu_{n})\setminus Y\neq\emptyset$ and by invariance of $\operatorname{supp}(\mu_{n})$, also $\operatorname{supp}(\mu_{n})\cap K^{\prime}\neq\emptyset$. The Birkhoff Ergodic Theorem implies that $\operatorname{orb}_{f}(x)$ is dense in $\operatorname{supp}(\mu_{n})$ for $\mu_{n}$-a.e. $x$. By the Ye & Zhang result, see [39, Proposition 4.4], we can pick $x$ such that additionally $h_{\text{\tiny top}}(x)\geq h^{\prime}$. But there is $k\in{\mathbb{N}}$ such that $f^{k}(x)\in K^{\prime}$. But then there is a compact neighbourhood $K^{\prime\prime}\owns x$ such that $f^{k}(K^{\prime\prime})\subset K$, which implies that $h_{\text{\tiny top}}(x)\leq h_{\text{\tiny top}}(z)<h^{\prime}$. This contradiction proves the theorem. $\Box$
@@
The assumption that the measures of maximal entropy are fully supported is not entirely automatic, although it is conjectured that on manifolds the measure of maximal entropy (if existent) is automatically fully supported if the map is topologically transitive, see [40, 14] for results supporting this conjecture. In the context of subshifts $(X,\sigma)$ of positive entropy (so $X$ is a Cantor set, not a manifold), Kwietniak et al. [25] presented a counter-example with all of its measures of maximal entropy supported on a proper subsets of $X$. Loosely inspired by their construction, we have the following topologically transitive (and in fact coded) subshift with a non-constant entropy function.
@@
We construct a coded subshift of $\{0,1,2\}^{\mathbb{Z}}$ on which the local entropy function is not constant. Let $\{w_{k}\}_{k\in{\mathbb{N}}}$ be a denumeration of all finite words in $\{0,1\}^{+}$ such that $|w_{k}|\leq|w_{k+1}|$ for all $k\in{\mathbb{N}}$. Let ${\mathcal{C}}=\{C_{k}\}_{k\in{\mathbb{N}}}$ the collection of code words, where
@@
A coded shift is the shift of which the language consists of subwords of free concatenations of code words, see [26] and [31]. The structure of the code words in ${\mathcal{C}}$ implies that $22$ is a synchronizing word (i.e., if $u22$ and $22v$ are both allowed words in $X_{{\mathcal{C}}}$, then so is $u22v$), but more importantly, $h_{\text{\tiny top}}(X_{{\mathcal{C}}},\sigma)\geq\log 2$, because the full shift on $\{0,1\}$ is a subshift of $X_{{\mathcal{C}}}$.
Now consider the cylinder set $[2.2]$ and let $z\in[2.2]$ be such that $\operatorname{orb}_{\sigma}(z)$ is dense in $X_{{\mathcal{C}}}$ and $z\in K\subset[2.2]$ any closed neighbourhood. From the theory of coded shifts, see [12, Section 3.3], we can compute the exponential growth rate of the number of centered words $x$ of length $12n+$ 1 and with $x_{-1}=x_{0}=2$, namely, this is the unique positive solution $h$ of $\sum_{k}e^{-h|C_{k}|}=1$, and it is clear that $h<\log 2$. It follows that $h_{\text{\tiny top}}(z)\leq h<h_{\text{\tiny top}}(X_{{\mathcal{C}}},\sigma)$. In this case, work of Pavlov [31, Theorem 1.1] implies that its measures of maximal entropy are supported on $\overline{\operatorname{orb}_{\sigma}(z)}\setminus\operatorname{orb}_{\sigma}(z)$, i.e., $\{0,1\}^{\mathbb{Z}}$. On the other hand, for every closed neighbourhood $K$ of $z$, there is $n$ such that $\sigma^{n}(K)\supset[2.2]$, and therefore $h_{\text{\tiny top}}(z)=\log h<h_{\text{\tiny top}}(X_{{\mathcal{C}}},\sigma)=\log 2$.
Without the assumption of transitivity, maps with non-constant entropy are not hard to find. A simple example of a map where $h_{\text{\tiny top}}(z)$ is not constant, and in fact, $\sup_{z}h_{\text{\tiny top}}(z)=\infty$, is given in Figure 1; see [23] for Hölder continuous maps of this type. Maps of this type (i.e., with infinitely many transitive components, each with its own value for local entropy, and infinite entropy altogether, are $C^{0}$-generic on any manifold of dimension $\geq 2$, see [17]. For such examples, the values that local entropy assume forms a countable dense subset of $[0,\infty)$, and with some more work (but giving up on genericity) one can find smooth examples with these properties.

### v3

@@ Abstract
In 2007, Ye & Zhang introduced a version of local topological entropy. Since their entropy function is, as we show under mild conditions, constant for topologically transitive dynamical systems, we propose to adjust the notion in a way that does not neglect the initial transient part of an orbit. We investigate the properties of this “transient” version, which we call translocal entropy, and compute it in terms of Lyapunov exponents for various dynamical systems. We also investigate how this adjustment affects measure-theoretic local (Brin-Katok) entropy and local pressure functions, generalizing some partial variational principles of Ma & Wen.
Mathematics Subject Classification (2010): Primary: 37B40, Secondary: 37A35.
Keywords. topological entropy, local entropy, Brin-Katok entropy, variational principle, topological pressure, local measure-theoretic pressure.
@@ 1 Introduction and motivation
Topological entropy is related to measure-theoretic entropy via the variational principle: If $f$ is a continuous map of a compact metric space $(X,d)$ (an assumption we will make throughout the paper), then $h_{\text{\tiny top}}(f)=\sup_{\mu}h_{\mu}(f)$, where the supremum is taken over all $f$-invariant probability measures $\mu$ on $X$. Measures $\mu$ such that $h_{\text{\tiny top}}(f)=h_{\mu}(f)$ are called measures of maximal entropy or maximal measures. In thermodynamic formalism, one tries to maximize the sum of entropy and energy (represented by an integral over a potential function $\phi:X\to{\mathbb{R}}$), called free energy or measure-theoretic pressure. The topological pressure can then be defined by the variational principle as $P_{top}(f)=\sup_{\mu}\{h_{\mu}+\int_{X}\phi\,d\mu\}$, see e.g. [39, 45].
Local versions of topological entropy were used by Bowen [10] to get an upper estimate of the difference between $h_{\mu}(f)$ and the entropy of a partition with diameter less than $\varepsilon$ and by Misiurewicz [34], who proved that vanishing of his local entropy implies existence of a maximal measure. Different versions were used by Newhouse [35], and by Buzzi & Ruette [15, 16], to address questions about the (unique) existence of a measure of maximal entropy. All these notions rely on Bowen balls (also called dynamical balls):
@@
A local view on measure-theoretic entropy goes back to at least Brin & Katok [12]. In particular, Brin-Katok local entropy $h_{\mu}(f,x)$ (see (7) in Section 3.1 for the definition), when integrated over the space yields the global measure-theoretic entropy:
$\int h_{\mu}(f,x)\,d\mu(x)=h_{\mu}(f).$
@@
In 2007, Ye & Zhang [49] introduced a new version of local entropy, which they called entropy function $h_{\text{\tiny top}}:X\to[0,\infty]$.
Their definition, see (3) below, is the starting point of our paper, but one soon discovers that under mild conditions the entropy function is constant: for any $z\in X$ one has $h_{\text{\tiny top}}(z)\equiv h_{\text{\tiny top}}(f)$ for very natural dynamical systems, see Corollary 2.1. It is not straightforward to find maps with non-constant entropy functions, see Section 2.1. The reason is that in the Ye & Zhang approach, only the limit behaviour of the orbit of $z$ counts, not the “transient” behaviour at finite time scales. The entropy function does not see if entropy is created “immediately” in small neighbourhoods of $z$, or only at very distant time scales.
For this reason, we propose an adjustment to Ye & Zhang’s definition, which we call the translocal entropy function and denote it as $h_{\omega}$, that takes into account the time scale needed for Bowen balls to reach unit size. As a result, $h_{\omega}(z)$ depends non-trivially on $z\in X$, and Lyapunov exponents at $z$ play a central role. This interaction is most transparent in smooth one-dimensional maps, see Theorem 2.1, where each point has at most one Lyapunov exponent (because there is only one direction in which derivatives can be taken). As representatives of higher-dimensional maps, we study toral automorphisms $F_{A}:{\mathbb{T}}^{d}\to{\mathbb{T}}^{d}$ based on a $d\times d$ matrix $A$. Theorem 2.2 shows that its translocal entropy function is determined by the Lyapunov exponents, i.e., the logarithms of the eigenvalues of $A$.
@@ 1.2 Structure of the paper
Our paper is organized as follows. In Section 2, we introduce the entropy function by Ye & Zhang, and study when it is constant and when not. In Section 2.2, we give our adaptation of the entropy function, and give its basic and more advanced properties in the remaining subsections of Section 2.
@@ 2 Local topological entropy in sense of Ye & Zhang
Let $f:X\to X$ be a continuous map on a compact metric space $(X,d)$. A set $A$ is called $(n,\varepsilon)$-separated if $d(f^{j}(x),f^{j}(y))>\varepsilon$ for some $0\leq j<n$ and all distinct $x,y\in A$. For a closed subset $K\subset X$, we define the topological entropy [^1] This is in line with the Bowen-Dinaburg definition [9, 18], used in e.g. [38, 41, 42], who were eventually interested in the multifractal properties of a measure-theoretic version of local entropy, in the line of the Brin-Katok entropy which we discuss in Section 3.1. $h_{\text{\tiny top}}(f,K)$ restricted to $K$ as the limit of exponential growth rates of $(n,\varepsilon)$-separated subsets of $K$ as $\varepsilon\to 0$, that is
@@
where $S(n,\varepsilon,K)$ is the maximal cardinality of $(n,\varepsilon)$-separated subsets of $K$.
Can one modify (1) into a pointwise definition of topological entropy by shrinking $K$ to a point? One way of doing this is
$\lim_{\delta\to 0}\inf_{K\in{\mathcal{K}}_{\delta}(x)}h_{\text{\tiny top}}(f,K),$ (2)
where ${\mathcal{K}}_{\delta}(x)$ is the collection of closed neighborhoods of $x$ of diameter $\leq\delta$. A point $x\in X$ is an entropy point if there are arbitrarily small closed neighbourhoods $K$ such that $h_{\text{\tiny top}}(f,K)>0$ and a full entropy point if in addition $h_{\text{\tiny top}}(f,K)=h_{\text{\tiny top}}(f)$, see [49, Section 2]. Hence, only maps with positive entropy can have full entropy points.
A map $f:X\to X$ is topologically exact or locally eventually onto if for every open $U\subset X$, there is $N\in{\mathbb{N}}$ such that $f^{N}(U)=X$. However, if $f$ is locally eventually onto, then there is $N_{K}$ such that $f^{N_{K}}(K)=X$. Notice that $(n-N_{k},\varepsilon)$-separated points $y_{1},y_{2}\in f^{N_{K}}(K)$ determine points $x_{1},x_{2}\in K$, with $y_{1}=f^{N_{K}}(x_{1})$ and $y_{2}=f^{N_{K}}(x_{2})$, such that $x_{1},x_{2}$ are $(n,\varepsilon)$-separated. Therefore $S(n,\varepsilon,K)\geq S(n-N_{k},\varepsilon,f^{N_{K}}(K))$. Thus the quantity in (2) is constant and equal to $h_{\text{\tiny top}}(f)$.
Instead, in [49, Definition 4.1] the local entropy function is defined as
$h_{\text{\tiny top}}(x)=\lim_{\varepsilon\to 0}h_{\text{\tiny top}}(x,\varepsilon),\qquad\text{for~~~}h_{\text{\tiny top}}(x,\varepsilon)=\lim_{\delta\to 0}\inf_{K_{x}\in{\mathcal{K}}_{\delta}(x)}\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K_{x}).$ (3)
This differs from (2) in the order in which the limits $\varepsilon\to 0$ and $\delta\to 0$ are taken. In general, $h_{\text{\tiny top}}(x)=h_{\text{\tiny top}}(f(x))$, for any point $x\in X$, so the local entropy is constant along forward orbits, see [49, Proposition 4.4(1)]. In fact, we have the following.
@@
Proof. Fix $\delta>0.$ Choose a point $x\in X$ with a dense orbit and a point $y\in X$ with $K_{y}\in{\mathcal{K}}_{\delta}(y).$ Then, there exists $n_{0}\geq 0$ such that $f^{n_{0}}(x)\in K_{y}$ and a set $K_{x}\in{\mathcal{K}}_{\delta}(x)$ such that $f^{n_{0}}(K_{x})\subset K_{y}$. For any $(n,\varepsilon)$-separated points $y_{1},y_{2}\in f^{n_{0}}(K_{x})$ there exist $(n+n_{0},\varepsilon)$-separated points $x_{1},x_{2}\in K_{x}$ (i.e. $y_{1}=f^{n_{0}}(x_{1})$ and $y_{2}=f^{n_{0}}(x_{2})$). Thus,
$S(n+n_{0},\varepsilon,K_{x})\leq S(n,\varepsilon,f^{n_{0}}(K_{x}))\leq S(n,\varepsilon,K_{y})$
and taking limsup with $n\to\infty$ we get
$\limsup_{n\to\infty}\frac{n+n_{0}}{n}\cdot\frac{1}{n+n_{0}}\log S(n+n_{0},\varepsilon,K_{x})\leq\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K_{y}).$
Applying (3) and taking the infimum over $K\in{\mathcal{K}}_{\delta}$ on both sides, we obtain
$h_{\text{\tiny top}}(x,\varepsilon)=\inf_{K_{x}\in{\mathcal{K}}_{\delta}(x)}\limsup_{n\to\infty}\frac{1}{n}\log S(n+n_{0},\varepsilon,K_{x})\leq\inf_{K_{y}\in{\mathcal{K}}_{\delta}(y)}\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K_{y})=h_{\text{\tiny top}}(y,\varepsilon),$
and hence $h_{\text{\tiny top}}(x)\leq h_{\text{\tiny top}}(y)$. $\Box$
@@ 2.1 When is the entropy function non-constant?
One can argue the merits of local entropy function that is equal to $h_{\text{\tiny top}}(f)$ everywhere. In this case, the behavior near a single orbit suffices to estimate the topological entropy, but on the other hand, such an entropy function cannot indicate any inhomogeneity in where in the system, and how rapidly, chaos is created.
The next theorem shows that the Ye & Zhang local entropy function is indeed constant under mild assumptions. A version of this theorem, for much more specific dynamical systems, can be found in [8].
Topological exactness gives a simple sufficient condition for the entropy function to be constant.
@@
If $f:X\to X$ is topologically exact map, then $h_{\text{\tiny top}}(z)\equiv h_{\text{\tiny top}}(f)$ for any $z\in X$.
@@
Proof. Given a closed neighbourhood $K\owns z$, take $N$ such that $f^{N}(K)\supset f^{N}(\mathring{K})=X$, where $\mathring{K}$ indicates the interior of $K.$ Then $\frac{1}{n}\log S(n,\varepsilon,K)\geq\frac{1}{n}\log S(n-N,\varepsilon,X)=\frac{n-N}{n}\frac{1}{n-N}\log S(n-N,\varepsilon,X)$. So after taking limits $n\to\infty$ and $\varepsilon\to 0$, we find $h_{\text{\tiny top}}(z)\geq h_{\text{\tiny top}}(f)$. $\Box$
Ye & Zhang [49, Theorem 3.7] prove, based on a lemma by Katok, that for every ergodic $f$-invariant measure $\mu$:
$\lim_{\varepsilon\to 0}\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K):K\text{ is a Borel set with }\mu(K)>0\}\geq h_{\mu}(f).$ (4)
This allows us to derive the following corollary:
Corollary 2.1
@@
Let $f:X\to X$ be a continuous topological transitive map defined on a compact metric space $(X,d)$, and assume that it has at least one measure of maximal entropy. If $h_{\text{\tiny top}}(f)=0$ or if each measure of maximal entropy is fully supported, then $h_{\text{\tiny top}}(x)=h_{\text{\tiny top}}(f)$ for every $x\in X$.
Proof. If $h_{\text{\tiny top}}(f)=0,$ then of course the corollary is true. Since all measure of maximal entropy are fully supported, certainly there is a fully supported ergodic $f$-invariant measure $\mu$ of maximal entropy. Therefore $\mu(K)>0$ for every compact neighbourhood of every $z\in X$. Fix $\delta>0$ and $x\in X.$ Since
$\inf$ $\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K_{x}):K_{x}\in{\mathcal{K}}_{\delta}(x)~and~\mu(K_{x})>0\}$
$\geq\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K):K\text{ is a Borel set with }\mu(K)>0\},$
we have
$\lim_{\delta\to 0}$ $\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K_{x}):K_{x}\in{\mathcal{K}}_{\delta}(x)~and~\mu(K_{x})>0\}$
$\geq\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K):K\text{ is a Borel set with }\mu(K)>0\}.$
Taking limit with $\varepsilon\to 0$ we get
@@
$h_{\text{\tiny top}}(x)$ $=$ $\lim_{\varepsilon\to 0}\lim_{\delta\to 0}\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K_{x}):K_{x}\in{\mathcal{K}}_{\delta}(x)~and~\mu(K_{x})>0\}$
$\geq$ $\lim_{\varepsilon\to 0}\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K):K\text{ is a Borel set with }\mu(K)>0\}.$
Formula (4) and the assumption that $\mu$ is a measure of maximal entropy yield
$h_{\text{\tiny top}}(f)\geq h_{\text{\tiny top}}(x)\geq h_{\mu}(f)=h_{\text{\tiny top}}(f).$
$\Box$
@@
The assumption that the measures of maximal entropy are fully supported is not entirely automatic, although it is conjectured that on manifolds the measure of maximal entropy (if existent) is automatically fully supported if the map is topologically transitive, see [50] for results supporting this conjecture. In the context of subshifts $(X,\sigma)$ of positive entropy (so $X$ is a Cantor set, not a manifold), Kwietniak et al. [30] presented a counter-example with all of its measures of maximal entropy supported on a proper subsets of $X$. Loosely inspired by their construction, we have the following topologically transitive (and in fact coded) subshift with a non-constant entropy function.
@@
We construct a coded subshift of $\{0,1,2\}^{\mathbb{Z}}$ on which the local entropy function is not constant. Let $\{w_{k}\}_{k\in{\mathbb{N}}}$ be an enumeration of all words in $\bigcup_{n\geq 1}\{0,1\}^{n}$ such that $|w_{k}|\leq|w_{k+1}|$ for all $k\in{\mathbb{N}}$. Let ${\mathcal{C}}=\{C_{k}\}_{k\in{\mathbb{N}}}$ be the collection of code words, where
@@
A coded shift $X_{{\mathcal{C}}}$ is the shift of which the language consists of subwords of free concatenations of code words, see [31] and [36]. The structure of the code words in $\mathcal{C}$} implies that $22$ is a synchronizing word (i.e., if $u22$ and $22v$ are both allowed words in $X_{{\mathcal{C}}}$, then so is $u22v$), but more importantly, $h_{\text{\tiny top}}(X_{{\mathcal{C}}},\sigma)\geq\log 2$, because the full shift on $\{0,1\}$ is a subshift of $X_{{\mathcal{C}}}$.
Now consider the cylinder set $2.2]=\{z\in\{0,1,2\}^{{\mathbb{Z}}}:z_{-1}=z_{0}=2\}$ and let $z\in[2.2]$ be such that $\operatorname{orb}_{\sigma}(z)$ is dense in $X_{{\mathcal{C}}}$ and $K\subset[2.2]$ is a closed neighbourhood of $z$. From the theory of coded shifts, see [14, Section 3.3], we can compute the exponential growth rate of the number of centered words $x$ of length $2n+1$ and with $x_{-1}=x_{0}=2$. Namely, this is the unique positive solution $h$ of $\sum_{k}e^{-h|C_{k}|}=1$, and it is clear that $h<\log 2$. It follows that $h_{\text{\tiny top}}(z)\leq h<h_{\text{\tiny top}}(X_{{\mathcal{C}}},\sigma)$. In this case, work of Pavlov [36, Theorem 1.1] implies that its measures of maximal entropy are supported on $\overline{\operatorname{orb}_{\sigma}(z)}\setminus\operatorname{orb}_{\sigma}(z)$, i.e., $\{0,1\}^{\mathbb{Z}}$. On the other hand, for every closed neighbourhood $K$ of $z$, there is $n$ such that $\sigma^{n}(K)\supset[2.2]$, and therefore $h_{\text{\tiny top}}(z)=\log h<h_{\text{\tiny top}}(X_{{\mathcal{C}}},\sigma)=\log 2$.
Without the assumption of transitivity, maps with non-constant entropy are not hard to find. A simple example of a map where $h_{\text{\tiny top}}(z)$ is not constant, and in fact, $\sup_{z}h_{\text{\tiny top}}(z)=\infty$, is given in Figure 1, see [27] for Hölder continuous maps of this type. Maps of this type (i.e., with infinitely many transitive components, each with its own value for local entropy, and infinite entropy altogether), are $C^{0}$-generic on any manifold of dimension $\geq 2$, see [20]. For such examples, the values $E$ that local entropy assumes forms a countable subset of $0,\infty)$. With some more work (but giving up on genericity) one can find smooth examples where $E$ is dense in $0,\infty)$, as well.